WAEC Further Mathematics 2011 past questions and answers

58 questions. Try each one, then open the answer to check yourself and read the explanation.

Question 1

Binary Operations

A binary operation * is defined on the set of real numbers, R, by a * b = a + b − 1. Find the identity element under the operation *.

  1. A.−1
  2. B.0
  3. C.1
  4. D.2
Show the answer and explanation

Answer: C (1)

Explanation

If e is the identity, a * e = a + e − 1 = a, so e − 1 = 0 and e = 1.

Open this question

Question 2

Trigonometry

Express 75° in radians, leaving your answer in terms of π.

  1. A.5π/12
  2. B.3π/4
  3. C.5π/6
  4. D.7π/6
Show the answer and explanation

Answer: A (5π/12)

Explanation

180° = π radians, so 75° = 75π/180 = 5π/12.

Open this question

Question 3

Indices and Logarithms

If log93 + 2x = 1, find x.

  1. A.−1/2
  2. B.−1/4
  3. C.1/4
  4. D.1/2
Show the answer and explanation

Answer: C (1/4)

Explanation

log93 = ½ because 9½ = 3. So ½ + 2x = 1, 2x = ½ and x = ¼.

Open this question

Question 4

Trigonometry

Evaluate cos(π/2 + π/3)

  1. A.2/√3
  2. B.−√3/2
  3. C.√3/4
  4. D.4/√3
Show the answer and explanation

Answer: B (−√3/2)

Explanation

cos(π/2 + π/3) = cos(π/2)cos(π/3) − sin(π/2)sin(π/3) = 0 − √3/2 = −√3/2.

Open this question

Question 5

Polynomials

Find the remainder when 5x3 + 2x2 − 7x − 5 is divided by (x − 2).

  1. A.−15
  2. B.−23
  3. C.29
  4. D.49
Show the answer and explanation

Answer: C (29)

Explanation

By the remainder theorem the remainder is f(2) = 40 + 8 − 14 − 5 = 29.

Open this question

Question 6

Functions

A function is defined by f(x) = (3x + 1)/(x2 − 1). Find f(−3).

  1. A.−1¼
  2. B.−1
  3. C.4/5
  4. D.1
Show the answer and explanation

Answer: B (−1)

Explanation

f(−3) = (−9 + 1)/(9 − 1) = −8/8 = −1.

Open this question

Question 7

Indices and Logarithms

Simplify: ∛(8/27) − (4/9)−1/2

  1. A.−5/6
  2. B.−4/27
  3. C.0
  4. D.2/9
Show the answer and explanation

Answer: A (−5/6)

Explanation

∛(8/27) = 2/3 and (4/9)−1/2 = (9/4)1/2 = 3/2. So 2/3 − 3/2 = −5/6.

Open this question

Question 8

Polynomials

Solve 3x2 + 4x + 1 > 0.

  1. A.x < −1, x < −⅓
  2. B.x > −1
  3. C.x > −⅓, x < −1
  4. D.x < −⅓, x < −⅓
Show the answer and explanation

Answer: C (x > −⅓, x < −1)

Explanation

3x2 + 4x + 1 = (3x + 1)(x + 1). The product is positive when both factors have the same sign: x > −⅓ or x < −1.

Open this question

Question 9

Coordinate Geometry

The equation of a circle is 3x2 + 3y2 + 6x − 12y + 6 = 0. Find its radius.

  1. A.1
  2. B.√3
  3. C.√11
  4. D.6
Show the answer and explanation

Answer: B (√3)

Explanation

Divide by 3: x2 + y2 + 2x − 4y + 2 = 0, i.e. (x + 1)2 + (y − 2)2 = 1 + 4 − 2 = 3. The radius is √3.

Open this question

Question 10

Functions

f(x) = p + qx, where p and q are constants. If f(1) = 7 and f(5) = 19, find f(3).

  1. A.13
  2. B.15
  3. C.17
  4. D.26
Show the answer and explanation

Answer: A (13)

Explanation

p + q = 7 and p + 5q = 19 give 4q = 12, q = 3 and p = 4. So f(3) = 4 + 9 = 13.

Open this question

Question 11

Polynomials

The sum and product of the roots of a quadratic equation are 4/7 and 5/7 respectively. Find its equation.

  1. A.7x2 − 4x − 5 = 0
  2. B.7x2 − 4x + 5 = 0
  3. C.7x2 + 4x − 5 = 0
  4. D.7x2 + 4x + 5 = 0
Show the answer and explanation

Answer: B (7x2 − 4x + 5 = 0)

Explanation

The equation is x2 − (sum)x + product = 0: x2 − (4/7)x + 5/7 = 0. Multiply by 7: 7x2 − 4x + 5 = 0.

Open this question

Question 12

Differentiation

f(x) = (x2 + 3)2 is defined on the set of real numbers R. Find the gradient of f(x) at x = 1/2.

  1. A.4.0
  2. B.6.5
  3. C.5.0
  4. D.10.6
Show the answer and explanation

Answer: B (6.5)

Explanation

f'(x) = 2(x2 + 3) × 2x = 4x(x2 + 3). At x = ½: 2 × 3.25 = 6.5.

Open this question

Question 13

Differentiation

Find limx→3 (x + 3)/(x2 − x − 12)

  1. A.−1
  2. B.−1/7
  3. C.1/7
  4. D.1
Show the answer and explanation

Answer: A (−1)

Explanation

The denominator is not zero at x = 3, so substitute: (3 + 3)/(9 − 3 − 12) = 6/(−6) = −1.

Open this question

Question 14

Differentiation

If y2 + xy − x = 0, find dy/dx

  1. A.(1 − y)/2y
  2. B.(1 − 2y)/x
  3. C.(1 − y)/(x + 2y)
  4. D.1/(x + 2y)
Show the answer and explanation

Answer: C ((1 − y)/(x + 2y))

Explanation

Differentiate implicitly: 2y(dy/dx) + x(dy/dx) + y − 1 = 0, so (2y + x)(dy/dx) = 1 − y and dy/dx = (1 − y)/(x + 2y).

Open this question

Question 15

Coordinate Geometry

A line is perpendicular to 3x − y + 11 = 0 and passes through the point (1, −5). Find its equation.

  1. A.3y − x − 14 = 0
  2. B.3x + y + 1 = 0
  3. C.3y + x + 1 = 0
  4. D.3y + x + 14 = 0
Show the answer and explanation

Answer: D (3y + x + 14 = 0)

Explanation

The given line has gradient 3, so the perpendicular has gradient −⅓. y + 5 = −⅓(x − 1) gives 3y + 15 = −x + 1, i.e. 3y + x + 14 = 0.

Open this question

Question 16

Indices and Logarithms

Solve 92x+1 = 813x+2.

  1. A.−3/4
  2. B.−2/3
  3. C.¾
  4. D.3/2
Show the answer and explanation

Answer: A (−3/4)

Explanation

81 = 92, so 92x+1 = 96x+4. Then 2x + 1 = 6x + 4, −4x = 3 and x = −¾.

Open this question

Question 17

Functions

The inverse of a function is given by f−1 : x → (x + 1)/4. Find f.

  1. A.f : x → 4x − 1
  2. B.f : x → 4x + 1
  3. C.f : x → 4x − ½
  4. D.f : x → x − ¼
Show the answer and explanation

Answer: A (f : x → 4x − 1)

Explanation

Let y = (x + 1)/4; then x = 4y − 1. So f is the inverse of f−1: f(x) = 4x − 1.

Open this question

Question 18

Matrices and Determinants

If (3 2 / 7 x)(2 / 3) = (12 / 29), find the value of x. (The 2 × 2 matrix is written row by row; the others are column vectors written top / bottom.)

  1. A.5
  2. B.6
  3. C.7
  4. D.8
Show the answer and explanation

Answer: A (5)

Explanation

The second row gives 7 × 2 + 3x = 29, so 3x = 15 and x = 5.

Open this question

Question 19

Sequences and Series

The fourth term of a geometric sequence is 2 and the sixth term is 8. Find the common ratio.

  1. A.±1
  2. B.±2
  3. C.±3
  4. D.±4
Show the answer and explanation

Answer: B (±2)

Explanation

ar3 = 2 and ar5 = 8. Dividing, r2 = 4, so r = ±2.

Open this question

Question 20

Differentiation

What percentage increase in the radius of a sphere will cause its volume to increase by 45%?

  1. A.10%
  2. B.15%
  3. C.20%
  4. D.25%
Show the answer and explanation

Answer: B (15%)

Explanation

V = (4/3)πr3, so for small changes δV/V ≈ 3 δr/r. Then 45% = 3 × (δr/r), giving about 15%.

Open this question

Question 21

Surds

Evaluate 1/(1 − sin 60°) leaving your answer in surd form.

  1. A.1 − √3
  2. B.2 − √3
  3. C.4 − 2√3
  4. D.4 + 2√3
Show the answer and explanation

Answer: D (4 + 2√3)

Explanation

1/(1 − √3/2) = 2/(2 − √3). Multiply top and bottom by 2 + √3: 2(2 + √3)/(4 − 3) = 4 + 2√3.

Open this question

Question 22

Coordinate Geometry

Find the equation of a circle with centre (−3, −8) and radius 4√6.

  1. A.x2 − y2 − 6x + 16y + 23 = 0
  2. B.x2 + y2 + 6x + 16y − 23 = 0
  3. C.x2 + y2 + 6x − 16y + 23 = 0
  4. D.x2 + y2 + 6x + 16y + 23 = 0
Show the answer and explanation

Answer: B (x2 + y2 + 6x + 16y − 23 = 0)

Explanation

(x + 3)2 + (y + 8)2 = 96. Expanding: x2 + y2 + 6x + 16y + 73 − 96 = 0, i.e. x2 + y2 + 6x + 16y − 23 = 0.

Open this question

Question 23

Binomial Theorem

Determine the coefficient of x2 in the expansion of (a + 3x)6.

  1. A.18a2
  2. B.45a4
  3. C.135a4
  4. D.1215a2
Show the answer and explanation

Answer: C (135a4)

Explanation

The x2 term is 6C2a4(3x)2 = 15 × 9a4x2, so the coefficient is 135a4.

Open this question

Question 24

Statistics

The mean of 2, 5, (x + 2), 7 and 9 is 6. Find the median.

  1. A.5.5
  2. B.6.0
  3. C.6.5
  4. D.7.0
Show the answer and explanation

Answer: D (7.0)

Explanation

The sum is 5 × 6 = 30, so 25 + x = 30 and x = 5. The numbers are 2, 5, 7, 7, 9, whose middle value is 7.

Open this question

Question 25

Probability

The probability that Kofi and Ama hit a target in a shooting competition are 1/6 and 1/9 respectively. What is the probability that only one of them will hit the target?

  1. A.1/54
  2. B.13/54
  3. C.20/27
  4. D.41/64
Show the answer and explanation

Answer: B (13/54)

Explanation

P(only Kofi) + P(only Ama) = (1/6)(8/9) + (5/6)(1/9) = 8/54 + 5/54 = 13/54.

Open this question

Question 26

Permutations and Combinations

In how many ways can 3 perfects be chosen out of 8 perfects?

  1. A.6
  2. B.24
  3. C.56
  4. D.336
Show the answer and explanation

Answer: C (56)

Explanation

Order does not matter, so the number is 8C3 = (8 × 7 × 6)/(3 × 2 × 1) = 56.

Open this question

Question 27

Statistics

Find the standard deviation of the numbers 3, 6, 2, 1, 7 and 5.

  1. A.2.00
  2. B.2.16
  3. C.2.50
  4. D.2.56
Show the answer and explanation

Answer: B (2.16)

Explanation

Mean = 24/6 = 4. Squared deviations: 1, 4, 4, 9, 9, 1, total 28. s.d. = √(28/6) = √4.67 ≈ 2.16.

Open this question

Question 28

Statistics

The table shows the distribution of marks of students in a class. Find the upper class boundary of the modal class.
Marks: 5−7, 8−10, 11−13, 14−16, 17−19, 20−22
No. of students: 4, 7, 26, 41, 14, 8

  1. A.13
  2. B.16
  3. C.16.5
  4. D.22.5
Show the answer and explanation

Answer: C (16.5)

Explanation

The modal class is 14−16 (frequency 41). Its upper class boundary is 16.5.

Open this question

Question 29

Permutations and Combinations

If 3xC2 = 15, find the value of x.

  1. A.2
  2. B.4
  3. C.5
  4. D.6
Show the answer and explanation

Answer: A (2)

Explanation

3x(3x − 1)/2 = 15, so 9x2 − 3x − 30 = 0, i.e. 3x2 − x − 10 = (x − 2)(3x + 5) = 0. x must be a positive whole number, so x = 2.

Open this question

Question 30

Permutations and Combinations

Four doctors and two nurses are to sit round a circular table. In how many ways can this be done if the nurses are to sit together?

  1. A.48
  2. B.60
  3. C.240
  4. D.720
Show the answer and explanation

Answer: A (48)

Explanation

Tie the two nurses together as one unit: 5 units round a table can sit in 4! = 24 ways, and the nurses can swap places in 2 ways. 24 × 2 = 48.

Open this question

Question 31

Probability

A basket contains 3 red and 1 white identical balls. A ball is drawn from the basket at random. Calculate the probability that it is either white or red.

  1. A.⅓
  2. B.½
  3. C.¾
  4. D.1
Show the answer and explanation

Answer: D (1)

Explanation

Every ball is either white or red, so the probability is 1/4 + 3/4 = 1.

Open this question

Question 32

Dynamics

A force of 200 N acting on a body of mass 20kg initially at rest causes it to move a distance of 320m along a straight line for t seconds. Find the value of t.

  1. A.4s
  2. B.6s
  3. C.8s
  4. D.10s
Show the answer and explanation

Answer: C (8s)

Explanation

a = F/m = 200/20 = 10 ms−2. s = ½at2: 320 = 5t2, t2 = 64 and t = 8 s.

Open this question

Question 33

Statics

Two forces 10N and 15N act on an object at an angle of 120° to each other. Find the magnitude of the resultant force.

  1. A.5√5
  2. B.5√7
  3. C.7√5
  4. D.7√7
Show the answer and explanation

Answer: B (5√7)

Explanation

R2 = 102 + 152 + 2(10)(15)cos 120° = 100 + 225 − 150 = 175, so R = √175 = 5√7 N.

Open this question

Question 34

Dynamics

A body of mass 25kg changes its speed from 15ms−1 to 35ms−1 in 5 seconds by the action of an applied force F. Find the value of F.

  1. A.100 N
  2. B.375 N
  3. C.500N
  4. D.600 N
Show the answer and explanation

Answer: A (100 N)

Explanation

F = m(v − u)/t = 25 × (35 − 15)/5 = 25 × 4 = 100 N.

Open this question

Question 35

Dynamics

A particle starts from rest and moves in a straight line such that its velocity, V, at time t seconds is given by v = (3t2 − 2t)ms−1. Calculate the distance covered in the first 2 seconds.

  1. A.2m
  2. B.4m
  3. C.6m
  4. D.8m
Show the answer and explanation

Answer: B (4m)

Explanation

s = ∫v dt = t3 − t2 (s = 0 at t = 0). At t = 2: s = 8 − 4 = 4 m.

Open this question

Question 36

Dynamics

A particle starts from rest and moves in a straight line such that its velocity, V, at time t seconds is given by v = (3t2 − 2t)ms−1. Determine the acceleration when t = 2 seconds.

  1. A.4 ms−2
  2. B.6 ms−2
  3. C.8 ms−2
  4. D.10 ms−2
Show the answer and explanation

Answer: D (10 ms−2)

Explanation

a = dv/dt = 6t − 2. At t = 2: a = 12 − 2 = 10 ms−2.

Open this question

Question 37

Vectors

Given that q = 9i + 6j and r = 4i − 6j, which of the following statements is true?

  1. A.r and q are collinear
  2. B.r and q are perpendicular
  3. C.The magnitude of r is 52 units
  4. D.The projection of r on q is √117 units
Show the answer and explanation

Answer: B (r and q are perpendicular)

Explanation

q · r = 9 × 4 + 6 × (−6) = 36 − 36 = 0. A zero dot product means the vectors are perpendicular.

Open this question

Question 38

Functions

The functions f and g are defined on the set, R, of real numbers by f : x → x2 − x − 6 and g : x → x − 1. Find f ∘ g(3).

  1. A.−8
  2. B.−6
  3. C.−4
  4. D.−3
Show the answer and explanation

Answer: C (−4)

Explanation

g(3) = 2, then f(2) = 4 − 2 − 6 = −4.

Open this question

Question 39

Vectors

Find the unit vector in the direction of −5i + 12j.

  1. A.1/13(−5i − 12j)
  2. B.1/13(5i + 12j)
  3. C.1/13(−5i + 12j)
  4. D.1/13(5i + 12j)
Show the answer and explanation

Answer: C (1/13(−5i + 12j))

Explanation

|−5i + 12j| = √(25 + 144) = 13, so the unit vector is (1/13)(−5i + 12j).

Open this question

Question 40

Vectors

Find, correct to two decimal places, the acute angle between p = (13 / 14) and q = (12 / 5). (Column vectors written top / bottom.)

  1. A.23.52°
  2. B.24.50°
  3. C.29.52°
  4. D.29.82°
Show the answer and explanation

Answer: B (24.50°)

Explanation

p · q = 156 + 70 = 226, |p| = √365, |q| = 13. cos θ = 226/(13√365) ≈ 0.9099, so θ ≈ 24.50°.

Open this question

Theory question 101

Polynomials

If f(x) = 6x3 + 13x2 + 2x − 5 and f(−1) = 0, find the factors of f(x).

Show the model answer
Model answer

f(−1) = 0, so (x + 1) is a factor
Divide: f(x) = (x + 1)(6x2 + 7x − 5)
6x2 + 7x − 5 = (2x − 1)(3x + 5)
Factors: (x + 1), (2x − 1) and (3x + 5)

Explanation

Use the given root to find one factor, then factorise the quadratic quotient.

Open this question

Theory question 102

Polynomials

If α and β are the roots of the equation 2x2 − 7x + 4 = 0, find the equation whose roots are α/β and β/α.

Show the model answer
Model answer

α + β = 7/2, αβ = 2
Sum of new roots = (α2 + β2)/αβ = [(α + β)2 − 2αβ]/αβ = (49/4 − 4)/2 = 33/8
Product of new roots = (α/β)(β/α) = 1
Equation: x2 − (33/8)x + 1 = 0, i.e. 8x2 − 33x + 8 = 0

Explanation

Write α² + β² as (α + β)² − 2αβ.

Open this question

Theory question 103

Binomial Theorem

(a) Write down the binomial expansion of (2 − x)5 in ascending powers of x.
(b) Use your expansion in 3(a) to evaluate (1.98)5 correct to four decimal places.

Show the model answer
Model answer

(a) (2 − x)5 = 32 − 80x + 80x2 − 40x3 + 10x4 − x5
(b) 1.98 = 2 − 0.02, so x = 0.02
32 − 1.6 + 0.032 − 0.00032 + 0.0000016 − ... = 30.43168
(1.98)5 ≈ 30.4317 (4 d.p.)

Explanation

Choose x so that 2 − x = 1.98, i.e. x = 0.02 (not 0.2).

Open this question

Theory question 104

Indices and Logarithms

Solve: 2(2y+2) − 9(2y) = −2

Show the model answer
Model answer

22y+2 = 4(2y)2; let p = 2y
4p2 − 9p + 2 = 0
(4p − 1)(p − 2) = 0, p = ¼ or p = 2
2y = 2 gives y = 1; 2y = ¼ = 2−2 gives y = −2
y = 1 or y = −2

Explanation

Substitute p = 2^y to get a quadratic.

Open this question

Theory question 105

Vectors

(a) Write the following as column vectors:
r = (10 N, 090°)
q = (8 N, 135°)
(b) use your answer in 5(a) to find (r + q)

Show the model answer
Model answer

Using bearings, a force (F, θ) = (F sin θ / F cos θ) (east component / north component)
(a) r = (10 sin 90° / 10 cos 90°) = (10 / 0)
q = (8 sin 135° / 8 cos 135°) = (4√2 / −4√2) ≈ (5.66 / −5.66)
(b) r + q = (10 + 4√2 / −4√2) ≈ (15.66 / −5.66)

Explanation

Resolve each force into east and north components using its bearing; then add component by component.

Open this question

Theory question 106

Permutations and Combinations

(a) Evaluate 9P3/15C3 + 5C3/3P2 correct to two decimal places.
(b) A committee of 2 tutors and 5 pupils is to be formed among 6 tutors and 10 pupils. In how many ways can this be done, if one particular tutor must be on the committee and two particular pupils must not be on the committee?

Show the model answer
Model answer

(a) 9P3 = 504, 15C3 = 455, 5C3 = 10, 3P2 = 6
504/455 + 10/6 = 1.1077 + 1.6667 = 2.77 (2 d.p.)
(b) The particular tutor is in; choose 1 more tutor from the other 5: 5 ways
Leave out the 2 particular pupils; choose 5 pupils from 8: 8C5 = 56 ways
Number of committees = 5 × 56 = 280

Explanation

Remove the people whose places are already decided, then choose from those left.

Open this question

Theory question 107

Statistics

A tyre manufacturing company researched into the life span of one type of their moto-cycle tyres. The results were as follows:
Distance (100km): 10 − 19, 20 − 29, 30 − 39, 40 − 49, 50 − 59, 60 − 69
Number of tyres: 30, 69, 93, 57, 36, 15
(a) Draw a histogram for the distribution.
(b) Use the histogram to estimate the mode.

Show the model answer
Model answer

(a) Class boundaries 9.5, 19.5, 29.5, 39.5, 49.5, 59.5, 69.5 on the horizontal axis; bars of equal width 10 and heights 30, 69, 93, 57, 36, 15
(b) Modal class 29.5 − 39.5 (tallest bar); join the corners of the tallest bar to the adjacent bars and read where the lines cross
Mode ≈ 29.5 + (93 − 69)/[(93 − 69) + (93 − 57)] × 10 = 29.5 + 4 = 33.5, i.e. about 3350 km

Explanation

Use class boundaries so the bars touch; the mode is read from the crossing lines on the tallest bar.

Open this question

Theory question 108

Dynamics

A particle of mass 400g is moving under the action of two forces; F1 = (35N, 210°) F2 = (35√3N, 300°) and a resistance of 40N. Find the magnitude of the:
(a) resultant of F1 and F2;
(b) resultant force acting on the particle.

Show the model answer
Model answer

Components (east, north) from bearings: F1 = (35 sin 210°, 35 cos 210°) = (−17.5, −30.31)
F2 = (35√3 sin 300°, 35√3 cos 300°) = (−52.5, 30.31)
(a) F1 + F2 = (−70, 0); magnitude 70 N (due west). Check: the forces are at 90°, so √(352 + 3 × 352) = 70 N
(b) The resistance of 40 N opposes the motion: resultant = 70 − 40 = 30 N

Explanation

The bearings differ by 90°, so Pythagoras gives the resultant quickly.

Open this question

Theory question 109

Differentiation

(a) Find, from first principles, the derivative of f(x) = (x + 3)2
(b) Evaluate ∫12 (x + 1)(x2 − 2x + 2)/x2 dx

Show the model answer
Model answer

(a) f(x + h) − f(x) = (x + h + 3)2 − (x + 3)2 = 2h(x + 3) + h2
[f(x + h) − f(x)]/h = 2(x + 3) + h
As h → 0: f'(x) = 2(x + 3) = 2x + 6
(b) (x + 1)(x2 − 2x + 2) = x3 − x2 + 2, so the integrand is x − 1 + 2x−2
∫ = [x2/2 − x − 2/x] from 1 to 2 = (2 − 2 − 1) − (½ − 1 − 2) = −1 + 2.5 = 3/2

Explanation

For (b) multiply out the top first, then divide each term by x².

Open this question

Theory question 110

Matrices and Determinants

(a) If A = (−2 5 / 4 3) and B = (3 1 / 2 3) find the value of x and y such that BA = 2(3 7 / −2 x) + (y 4 / 12 −3). (Matrices are written row by row.)
(b) Two functions, f and g are defined by f : x → ½x + 1 and g : x → (5x − 1)/3. Find (i) g−1 (ii) g−1 ∘ f.

Show the model answer
Model answer

(a) BA = (3(−2) + 1(4) 3(5) + 1(3) / 2(−2) + 3(4) 2(5) + 3(3)) = (−2 18 / 8 19)
Right side = (6 + y 18 / 8 2x − 3)
6 + y = −2, so y = −8; 2x − 3 = 19, so x = 11
(b)(i) y = (5x − 1)/3 gives x = (3y + 1)/5, so g−1(x) = (3x + 1)/5
(ii) g−1(f(x)) = [3(½x + 1) + 1]/5 = (3x + 8)/10

Explanation

Multiply B by A (rows of B into columns of A), then compare corresponding entries.

Open this question

Theory question 111

Matrices and Determinants

The images of points (2, 3) and (4, 5) under a linear transformation A are (3, 4) and (5, 6) respectively. Find the:
(a) matrix A;
(b) inverse of A;
(c) point whose image is (−1, 1).

Show the model answer
Model answer

(a) Let A = (a b / c d) (row by row): 2a + 3b = 3, 4a + 5b = 5 give b = 1, a = 0
2c + 3d = 4, 4c + 5d = 6 give d = 2, c = −1
A = (0 1 / −1 2)
(b) |A| = 0 × 2 − 1 × (−1) = 1; A−1 = (2 −1 / 1 0)
(c) A−1(−1 / 1) = (2(−1) − 1(1) / 1(−1) + 0) = (−3 / −1); the point is (−3, −1)

Explanation

Set up two pairs of simultaneous equations from the images; use the inverse to go backwards.

Open this question

Theory question 112

Differentiation

(a) Find the equation of the tangent to curve x2/4 + y2 = 1 at point (1, √3/2).
(b) Express (3x + 2)/(x2 + x − 2) in partial fractions.

Show the model answer
Model answer

(a) x/2 + 2y(dy/dx) = 0, so dy/dx = −x/(4y)
At (1, √3/2): dy/dx = −1/(2√3) = −√3/6
y − √3/2 = −(√3/6)(x − 1)
√3x + 6y − 4√3 = 0 (or x + 2√3y − 4 = 0)
(b) x2 + x − 2 = (x − 1)(x + 2); 3x + 2 = A(x + 2) + B(x − 1)
x = 1: 5 = 3A, A = 5/3; x = −2: −4 = −3B, B = 4/3
(3x + 2)/(x2 + x − 2) = 5/[3(x − 1)] + 4/[3(x + 2)]

Explanation

Differentiate implicitly for the gradient; factorise the denominator for the partial fractions.

Open this question

Theory question 113

Probability

A survey conducted revealed that four out of every twenty taxi drivers do not have valid driving license. If 6 drivers are selected at random, calculate, correct to three decimal places, the probability that: (a) exactly 2; (b) more than 3; (c) at least 5 have valid driving license.

Show the model answer
Model answer

P(valid) = 16/20 = 0.8, P(not valid) = 0.2; binomial with n = 6
(a) 6C2(0.8)2(0.2)4 = 15 × 0.64 × 0.0016 = 0.015
(b) P(4) + P(5) + P(6) = 0.24576 + 0.393216 + 0.262144 = 0.901
(c) P(5) + P(6) = 0.393216 + 0.262144 = 0.655

Explanation

Four out of twenty have no licence, so the probability of a valid licence is 0.8.

Open this question

Theory question 114

Statistics

The table shows the frequency distribution of marks scored by some candidates in an examination.
Marks: 0 − 9, 10 − 19, 20 − 29, 30 − 39, 40 − 49, 50 − 59, 60 − 69, 70 − 79, 80 − 89, 90 − 99
Frequency: 2, 5, 8, 18, 20, 15, 5, 4, 2, 1
(a) Draw the cumulative frequency curve of the distribution.
(b) Use your graph to estimate the: (i) Semi − interquartile range of the distribution; (ii) percentage of candidates who passed with distinction if the least mark for distinction was 72.

Show the model answer
Model answer

(a) Upper class boundaries 9.5, 19.5, ..., 99.5 against cumulative frequencies 2, 7, 15, 33, 53, 68, 73, 77, 79, 80; join with a smooth curve
(b)(i) N = 80: Q1 = 20th mark ≈ 29.5 + (5/18) × 10 ≈ 32.3; Q3 = 60th mark ≈ 49.5 + (7/15) × 10 ≈ 54.2
Semi-interquartile range = (54.2 − 32.3)/2 ≈ 10.9 (accept about 10.5 to 11 from the graph)
(ii) Read the cumulative frequency at 71.5: about 73.8, so about 80 − 73.8 ≈ 6 candidates scored 72 or more
Percentage ≈ 6.2/80 × 100 ≈ 7.8% (about 8%)

Explanation

Quartiles are at N/4 and 3N/4 on the cumulative frequency axis.

Open this question

Theory question 115

Probability

A bag contains 4 red, 6 blue and 8 green identical marbles.
(a) If three marbles are drawn at random, without replacement, calculate the probability that: (i) all will be green; (ii) all will have the same colour.
(b) If each marble is replaced before another is drawn, calculate the probability that all will have the same colour.

Show the model answer
Model answer

Total 18 marbles
(a)(i) (8/18)(7/17)(6/16) = 336/4896 = 7/102
(ii) P(GGG) + P(BBB) + P(RRR) = 7/102 + (6 × 5 × 4)/4896 + (4 × 3 × 2)/4896 = (336 + 120 + 24)/4896 = 480/4896 = 5/51
(b) (8/18)3 + (6/18)3 + (4/18)3 = (512 + 216 + 64)/5832 = 792/5832 = 11/81

Explanation

Without replacement the numbers on top and bottom both drop by one each draw.

Open this question

Theory question 116

Statics

(a) An object P of mass 6.5kg is suspended by two light inextensible strings AP and BP. The strings make angles 50° and 60° respectively with the downward vertical. (i) Express the forces acting on P in component form. (ii) If P is at rest, write down the vector equation connecting all the forces. (iii) Calculate, correct to one decimal place, the tensions in the strings.
(b) A particle of mass 5kg moves with initial velocity (−1 / 2)ms−1 and final velocity (3 / 4)ms−1. Find the magnitude of its change in momentum. (Column vectors are written top / bottom.)

Show the model answer
Model answer

(a) Take g = 10 ms−2: weight W = 65 N. Let T1 (in AP) and T2 (in BP)
(i) T1 = (−T1 sin 50° / T1 cos 50°), T2 = (T2 sin 60° / T2 cos 60°), W = (0 / −65)
(ii) T1 + T2 + W = 0
(iii) Horizontal: T1 sin 50° = T2 sin 60°; vertical: T1 cos 50° + T2 cos 60° = 65
Solving (or Lami's theorem, angle between strings 110°): T1 = 65 sin 60°/sin 110° ≈ 59.9 N, T2 = 65 sin 50°/sin 110° ≈ 53.0 N
(with g = 9.8: 58.7 N and 51.9 N)
(b) Change in momentum = m(v − u) = 5[(3 / 4) − (−1 / 2)] = 5(4 / 2) = (20 / 10) Ns
Magnitude = √(400 + 100) = √500 = 10√5 ≈ 22.4 Ns

Explanation

Resolve horizontally and vertically; the horizontal parts of the two tensions must balance.

Open this question

Theory question 117

Vectors

(a) m(2 / 1) + n(−1 / 2) = (5 / −1) where m and n are scalars. Find the value of (m + n). (Column vectors are written top / bottom.)
(b) A(−1, 3), B(2, −1) and C(5, 3) are the vertices of ABC. (i) Express in column notation, the unit vectors parallel to AB and AC. (ii) Use a dot product to calculate BÂC, correct to the nearest degree.

Show the model answer
Model answer

(a) 2m − n = 5 and m + 2n = −1
Double the first and add: 5m = 9, m = 9/5; n = 2m − 5 = −7/5
m + n = 2/5
(b)(i) AB = (3 / −4), |AB| = 5, unit vector (3/5 / −4/5)
AC = (6 / 0), |AC| = 6, unit vector (1 / 0)
(ii) cos BÂC = (AB · AC)/(|AB||AC|) = 18/30 = 0.6
BÂC ≈ 53°

Explanation

Use AB = B − A and AC = C − A, not the position vectors themselves.

Open this question

Theory question 118

Dynamics

(a) Forces F1 = (2N, 210°), F2 = (4N, 120°) act on a particle of mass 7kg which is at rest. Calculate the: (i) acceleration of the particle; (ii) velocity of the particle after 3 seconds
(b) F1 = (2i + 3j)N, F2 = (−5j)N and F3 = (6i − 4j)N act on a body. Find the magnitude and direction of the fourth force that will keep the body in equilibrium.

Show the model answer
Model answer

(a) Bearings 210° and 120° differ by 90°, so R = √(22 + 42) = √20 = 2√5 ≈ 4.47 N
(components: F1 = (−1, −1.732), F2 = (3.464, −2), sum (2.464, −3.732))
(i) a = R/m = 4.47/7 ≈ 0.64 ms−2
(ii) v = u + at = 0 + 0.639 × 3 ≈ 1.92 ms−1
(b) F1 + F2 + F3 = 8i − 6j, so F4 = −8i + 6j
|F4| = √(64 + 36) = 10 N
Direction: tan−1(6/8) ≈ 36.9° above the negative x-axis, i.e. 143.1° to the positive x-axis (bearing ≈ 306.9°, N53.1°W)

Explanation

In equilibrium the fourth force is minus the sum of the other three.

Open this question